Apologies if nobody is present. Judith is OOTO and I am working on network equipment to bring up new purchases.
-David
GGUS Ticket #144542: Stage-in issues. Getting little from CERN regarding help debugging. Last update Judith removed the secondary lsm mover from the production queue, which was requested.
UC:
- New Workers are all up and added to the pool.
- A few new storage nodes added to pool, waiting on network equipment updates to finish the rest
- Working on getting network equipment online for the rest of the new purchases
IU
- New workers added to pool
- New SLATE node is set up
UIUC
- Purchases have been made. Still waiting on their arrival.